A sample of oxygen gas has a volume of 545 mL at 35 degree C. The...


7. A sample of oxygen gas has a volume of 545 mL at 35 degree C. The gas is heated to 151 degree C at constant pressure in a container that can contract or expand. What is the final volume of the oxygen gas?  A) 750 mL  B) 396 mL  C) 417 mL  D) 267 mL  E) 126 mL  


8. A 45 mL sample of nitrogen gas is cooled from 135 degree C to 15 degree C in a container that can be contract or expand at constant pressure, what is the new volume of the nitrogen gas? 

 A) 64 mL  B) 5.0 mL  C) 410 mL  D) 32mL  E) 41mL  


9. The temperature of a sample of argon in 365 at 740 middot mmHg and 25 degree C is lowered to 12 degree C. Assuming the volume of the container and the amount of gas is unchanged, calculate the new pressure of the argon.

  A) 0.468 atm  B) 0.931 atm  C) 1.02 atm  D) 1.54 atm  E) 2.03 atm  


10. 0.820 mole of hydrogen gas has a volume of 2.00 L at a certain temperature and pressure. What is the volume of 0.125 mol of this gas at the same temperature and pressure? 

 A) 0.0512L  B) 0.250L  C) 0.305L  D) 4.01L  E) 19.5L
Answer #1

Initial Final preas me constant Temp (2)= 151+273 Temp (i)=35+273 J2=424k Ti =308 K V2 = ? Volume (V)= 545 mL apply formula 0From 760 (121273) * 740 atm x 365 mL P2 X 365 mL (25+273) K preasure (B) = 0.931 atm Ans: ③ 0.931 atml volume () a moles m) 0